# A Moment-Conserving Discontinuous Galerkin Representation of the Relativistic Maxwellian Distribution

### v1

@@ 1 Introduction
The Maxwell-Jüttner distribution, originally described by Jüttner (1911), is the relativistic equivalent to the classical Maxwellian distribution. It is the local maximum entropy configuration of a relativistic system of particles and plays the same role in relativistic thermodynamics that the Maxwellian does in non-relativistic thermodynamics. These distributions are central to kinetic modeling, in which the particle distribution function evolves using the Boltzmann, or Vlasov, equation. In such calculations, the distribution function can be far from local thermodynamic equilibrium, and the Maxwell-Jüttner distribution is used for projecting initial conditions, computing approximate collision operators, and for computing differences from local thermodynamic equilibrium for studying and deriving reduced models.
@@ 2 Relativistic Moments and the Maxwell-Jüttner Distribution
The derivation of the Maxwell-Jüttner (MJ) in arbitrary dimensional form is presented in Chacon-Acosta et al. (2010). The result, presented in the fluid-stationary frame, is written in the normalized form for a $d$-dimensional momentum space as
$f^{MJ}(\mathbf{p})=\frac{n}{2c(2\pi)^{\frac{d-1}{2}}(m_{0}c)^{d}\theta^{\frac{d-1}{2}}K_{\frac{d+1}{2}}(1/\theta)}\exp{\left(\frac{-\gamma(\mathbf{p})}{\theta}\right)}.$ (1)
Where $\theta=kT/m_{0}c^{2}$, $\gamma(\mathbf{p})=\sqrt{1+\mathbf{p}\cdot\mathbf{p}/m_{0}^{2}c^{2}}$, and $K_{\nu}(1/\theta)$ is the modified Bessel function of the second kind evaluated at $1/\theta$. The distribution is described in terms of the fluid-stationary frame quantities, specifically the number density, $n$, and temperature, $T$. The momentum measured in this, the fluid-stationary frame, is $\mathbf{p}$. The remaining quantities have their usually assigned meanings: the speed of light, $c$, the Boltzmann constant $k$, and the particle rest mass, $m_{0}$.
The distribution function, equation 1, is invariant under Lorentz transformation. This allows us to write the distribution function in a new frame, observed from the simulation perspective, labeled the lab frame. Variables in the lab frame will be represented with primed quantities. Transforming the distribution to the lab frame requires Lorentz transforming unprimed to primed quantities. For the Maxwell-Jüttner distribution, the only quantity that requires transformation is the $\gamma(\mathbf{p})$ in the exponential.
@@
$f^{MJ}(\mathbf{p})=f^{\prime,MJ}(\mathbf{p^{\prime}})=\frac{n}{2c(2\pi)^{\frac{d-1}{2}}(m_{0}c)^{d}\theta^{\frac{d-1}{2}}K_{\frac{d+1}{2}}(1/\theta)}\exp{\left(\frac{-\Gamma(\gamma(\mathbf{p^{\prime}})-\mathbf{v}_{b}\cdot\mathbf{p}^{\prime}/(m_{0}c^{2}))}{\theta}\right)}.$ (2)
Here, $n$, and $T$ remain defined as fluid-stationary frame quantities for the density and the temperature. In addition, we define the Lorentz boost factor $\Gamma=1/\sqrt{1-v_{b}^{2}/c^{2}}$, henceforth called the gamma factor, between the fluid-stationary and the lab frame. Note the lab frame is the same as the simulation frame and these terms are often used interchangeably. In the transformation of $\gamma$,
$\gamma=\Gamma(\gamma^{\prime}-\mathbf{v}_{b}\cdot\mathbf{p}^{\prime}/(m_{0}c^{2})),$ (3)
we have suppressed the ($\mathbf{p}^{\prime}$) dependence for notational convenience because it is implied. This transform will be utilized again when computing stationary frame moments, $n$ and $T$, while the distribution function is in the lab frame.
To simplify these equations, we rescale the velocity, $\mathbf{v}$, by a factor $c$ to make it dimensionless. We also define the dimensionless quantity, $\mathbf{u}\equiv\gamma\mathbf{v}=\mathbf{p}/m_{0}$, which replaces all instances of $\mathbf{p}$. We will refer to $\mathbf{u}$ as the momentum, since it is $\mathbf{p}$ divided by a constant rest mass $m_{0}$ and also normalized by the speed of light, $c$. Additional normalization of the spatial lengths, $x$, to a code unit scale so $x$ is also dimensionless, and consequently the distribution function $f(\mathbf{u})$ is dimensionless as well.
The MJ distribution requires fluid stationary moments, $n$, $\mathbf{v}_{b}$, and $T$, to project the function onto the grid. These moments can either be specified, such as those given in initial conditions, or computed from an existing arbitrary distribution $f^{\prime}(\mathbf{u^{\prime}})$. In the latter case, we calculate the moments from within the lab frame where the distribution function exists. The next section provides the equations for computing the fluid stationary moments.
@@ 2.1 Moments
The most general form for computing the moments involves the calculation of the flux four-vector and the stress-energy tensor. From these we provide a procedure for both computing the needed MJ moments as well as other diagnostics such as transport coefficients. Measured in the lab frame for an arbitrary distribution, the flux four-vector and stress-energy tensor components are written, respectively, as [^1] For a comprehensive source on relativistic thermodynamics, see Vereshchagin & Aksenov (2017).
$F^{\prime\mu}=\int f^{\prime}(\mathbf{u^{\prime}})u^{\prime\mu}\frac{d^{3}u^{\prime}}{\gamma^{\prime}}$ (5)
$T^{\prime\mu\nu}=m_{0}\int f^{\prime}(\mathbf{u^{\prime}})u^{\prime\mu}u^{\prime\nu}\frac{d^{3}u^{\prime}}{\gamma^{\prime}}.$ (6)
Where the four vector $u^{\mu}$ is defined as $u^{\mu}\equiv(\gamma,u^{i})$.
To compute the boost between the lab and stationary frames we require the velocity $\mathbf{v}_{b}$. The velocity can be computed from the lab frame flux four-vector. Suggestively relabeling some quantities of the flux vector makes this more transparent. First, the component $F^{\prime\mu=0}$ is the density in the lab frame, $N=F^{\prime 0}$. Second, for clarity, the fluxes are relabeled $Nv_{b}^{i}=F^{\prime i}$ for $i=\{1,2,3\}$. It is clear then that the velocity is given by $\mathbf{v}_{b}=F^{\prime i}/F^{\prime 0}=N\mathbf{v}_{b}/N.$
The velocity, $\mathbf{v}_{b}$, gives us all the information we need to transform the lab frame quantities to the fluid stationary frame. Computing $\Gamma=1/\sqrt{1-v_{b}^{2}}$, where $v_{b}^{2}=\mathbf{v}_{b}\cdot\mathbf{v}_{b}$, the components of the Lorentz transform in Cartesian coordinates is given by
@@
We will call this specific Lorentz transform from the lab to the fluid frame $\Lambda$. The stationary frame flux and stress-energy tensor is computed by transforming the lab respective quantities,
$F^{\alpha}=\Lambda_{\phantom{\mu}\mu}^{\alpha}F^{\prime\mu}$ (8)
$T^{\alpha\beta}=\Lambda_{\phantom{\mu}\mu}^{\alpha}\Lambda_{\phantom{\nu}\nu}^{\beta}T^{\prime\mu\nu}.$ (9)
The transformed four-vector for the flux, $F^{\alpha}=(n,0,0,0)$, directly provides the fluid-stationary frame density $n$. Equipped with $T^{\alpha\beta}$, $F^{\prime\alpha}$, and $F^{\alpha}$, we can immediately compute the stationary frame density, velocity, and temperature,
$n=F^{0}=N/\Gamma$ (10)
$v_{b}^{i}=F^{\prime i}/F^{\prime 0}=Nv_{b}^{i}/N$ (11)
$P=nT=\frac{1}{3}(T^{11}+T^{22}+T^{33}).$ (12)
Since we only require the diagonal spatial components, we can simplify the expression for the pressure moment calculation to $\mu=\nu=1,2,3$ which yields (absorbing Boltzmann’s constant $k$ into $T$),
$P=nT=\frac{m_{0}}{d}\int f(\mathbf{u})u^{2}\frac{d^{3}u}{\gamma}.$ (13)
We can replace the fluid-stationary quantity, $u^{2}$, in the integrand with lab frame quantities, $u^{2}=\gamma^{2}-1=\Gamma^{2}(\gamma^{\prime}-\mathbf{v}_{b}\cdot\mathbf{u^{\prime}})^{2}-1$ from equation 3. In addition, the Lorentz transformation between the primed and unprimed volumes are
@@
The last line, equation 16, is a statement that the phase-space volume element is invariant. [^2] These results, pointed out by Zenitani (2015), are arrived at by considering a proper or canonical intervals of time, length, and momentum. For example, in the case of time, $d\tau$ = $dt/\gamma$ = $dt^{\prime}/\gamma^{\prime}$, where $\tau$ is the proper/canonical time interval. Therefore, transforming the volume element with equation 14, and using $f(\mathbf{u})=f^{\prime}(\mathbf{u^{\prime}})$, we have an expression for the fluid-stationary pressure measured in the lab frame
$P=nT=\frac{m_{0}}{d}\int f^{\prime}(\mathbf{u^{\prime}})(\Gamma^{2}(\gamma^{\prime}-\mathbf{v}_{b}\cdot\mathbf{u^{\prime}})^{2}-1)\frac{d^{3}u^{\prime}}{\gamma^{\prime}}.$ (17)
Dividing out the fluid-stationary frame number density, $n$, leaves us with the temperature, $T$, of the distribution in the fluid-stationary frame.
In summary, the necessary equations to calculate the moments begin with equations 5 and 11 to isolate $\mathbf{v}_{b}$. Then use $\mathbf{v}_{b}$ to compute $\Gamma=1/\sqrt{1-v_{b}^{2}}$ and then to compute $n$ via equation 10. Finally, $T$ is computed using equation 17. This provides all the moments of an arbitrary distribution needed to construct a Maxwell-Jüttner distribution with equivalent moments.
@@ 3.1 Division and Multiplication in DG, Computing $\Gamma$
Here we choose to consistently use a poly-order $p$ modal representation for the DG scheme. This is because in previous studies we have found that we can utilize modal representations and corresponding weak operations to minimize aliasing errors, which can greatly impact robustness in kinetic calculations. See Juno et al. (2018), Hakim & Juno (2020) and Hakim et al. (2020). We note that it may be favorable to use nodal representations and tolerate the aliasing errors for these computations because, even if they are more inaccurate, the procedure may be more robust to realizability issues.
The linear algebra calculations, which arise from weak equality of DG-represented values, can be solved exactly. Take, for example, the calculation of isolating $\mathbf{v}_{b}$ from $\langle N\mathbf{v}_{b}\rangle$. Here, the braces, delimited-⟨$\langle$..$\rangle$ … ⟩, denotes a combined quantity resulting from the moment computation. The expansion of the moments $N,\langle N\mathbf{v}_{b}\rangle$ in terms of an orthonormal set of basis functions in the configuration space, $\phi_{i}(\mathbf{x})$, where $\mathbf{x}$ is defined for an interval within a cell $x_{i}\in[-1,1]$ that has an associated volume, $I$. In each cell, DG quantities are represented as the sum of coefficients times the modal basis
@@
With these expressions, we can rewrite $N\mathbf{v}_{b}=\langle N\mathbf{v}_{b}\rangle$. Where the left-hand side is two separate quantities $N$ times $\mathbf{v}_{b}$ and the right-hand side is a single quantity $\langle N\mathbf{v}_{b}\rangle$. Integrating both sides of this equation gives a relation for the coefficients
[…]
Evaluating the integrals in parenthesis, the right hand side simplifies to the Kronecker delta, $\delta_{ik}$, for an orthonormal basis. The left hand side integral is a bit more involved, but can be calculated analytically before the simulation begins. The weak equality, equation 19 then can be written as
@@
[…]
@@
Hypothetically, if we desired to calculate $\mathbf{c}_{Nv_{b},k}$ given the other two coefficients, $c_{N,i}$, $\mathbf{c}_{v_{b},j}$ we could have simply stopped at equation 20.
@@
[…]

### v2

@@ 1 Introduction
The Maxwell-Jüttner distribution, originally described by Jüttner (1911), is the relativistic equivalent to the classical Maxwellian distribution. It is the local maximum entropy configuration of a relativistic system of particles and plays the same role in relativistic thermodynamics that the Maxwellian does in non-relativistic thermodynamics. These distributions are central to kinetic modeling, in which the particle distribution function evolves using the Boltzmann, or Vlasov, equation. In these calculations, while the distribution function can be far from local thermodynamic equilibrium, the Maxwell-Jüttner distribution is used for projecting initial conditions, computing approximate collision operators, and for computing differences from local thermodynamic equilibrium for studying and deriving reduced models.
@@ 2 Relativistic Moments and the Maxwell-Jüttner Distribution
The derivation of the Maxwell-Jüttner (MJ) in arbitrary dimensional form is presented in Chacon-Acosta et al. (2010). The result, presented in the fluid-stationary frame denoted with primed quantities, is written in the normalized form for a $d$-dimensional momentum space as
$f^{\prime,MJ}(\mathbf{p}^{\prime})=\frac{n}{2(2\pi)^{\frac{d-1}{2}}(m_{0})^{d}\theta^{\frac{d-1}{2}}K_{\frac{d+1}{2}}(1/\theta)}\exp{\left(\frac{-\gamma(\mathbf{p^{\prime}})}{\theta}\right)}.$ (1)
We adopt natural units $c=k=1$, and define $\theta=T/m_{0}$, $\gamma(\mathbf{p}^{\prime})=\sqrt{1+\mathbf{p}^{\prime}\cdot\mathbf{p}^{\prime}/m_{0}^{2}}$, and $K_{\nu}(1/\theta)$ as the modified Bessel function of the second kind evaluated at $1/\theta$. The distribution is described in terms of the fluid-stationary frame quantities, specifically the number density, $n$, and temperature, $T$. The momentum measured in this, the fluid-stationary frame, is $\mathbf{p}^{\prime}$. Finally, the particle rest mass, $m_{0}$, is left unspecified for simulations with multiple species.
The distribution function, equation (1), is invariant under Lorentz transformation. This allows us to write the distribution function in the frame observed from the simulation perspective, labeled the lab frame. Variables in the lab frame will be represented with unprimed quantities. Transforming the distribution to the lab frame requires Lorentz transforming unprimed to primed quantities. For the Maxwell-Jüttner distribution, the only quantity that requires transformation is the $\gamma(\mathbf{p}^{\prime})$ in the exponential.
@@
$f^{MJ}(\mathbf{p})=f^{\prime,MJ}(\mathbf{p^{\prime}})=\frac{n}{2(2\pi)^{\frac{d-1}{2}}(m_{0})^{d}\theta^{\frac{d-1}{2}}K_{\frac{d+1}{2}}(1/\theta)}\exp{\left(\frac{-\Gamma(\gamma(\mathbf{p})-\mathbf{v}_{b}\cdot\mathbf{p}/m_{0})}{\theta}\right)}.$ (2)
Here, $n$, and $T$ remain defined as fluid-stationary frame quantities for the density and the temperature. In addition, we define the Lorentz boost factor $\Gamma=1/\sqrt{1-v_{b}^{2}}$, henceforth called the gamma factor, between the fluid-stationary and the lab frame. Note the lab frame is the same as the simulation frame and these terms are often used interchangeably. In the transformation of $\gamma$,
$\gamma^{\prime}=\Gamma(\gamma-\mathbf{v}_{b}\cdot\mathbf{p}/m_{0}),$ (3)
we have suppressed the ($\mathbf{p}$) dependence for notational convenience because it is implied. This transform will be utilized again when computing fluid-stationary frame moments, $n$ and $T$, while the distribution function is in the lab frame.
To simplify these equations, we also define the dimensionless quantity, $\mathbf{u}\equiv\gamma\mathbf{v}=\mathbf{p}/m_{0}$, which replaces all instances of $\mathbf{p}$. We will refer to $\mathbf{u}$ as the momentum, since it is $\mathbf{p}$ divided by a constant rest mass $m_{0}$. Additional normalization of the spatial lengths, $x$, to a code unit scale so $x$ is also dimensionless, and consequently the distribution function $f(\mathbf{u})$ is dimensionless as well.
The MJ distribution requires fluid stationary moments, $n$, $\mathbf{v}_{b}$, and $T$, to project the function onto the grid. These moments can either be specified, such as those given in initial conditions, or computed from an existing arbitrary distribution $f(\mathbf{u})$. In the latter case, we calculate the moments from within the lab frame where the distribution function exists. The next section provides the equations for computing the fluid stationary moments.
@@ 2.1 Moments
The most general form for computing the moments involves the calculation of the flux four-vector and the stress-energy tensor. From these we provide a procedure for both computing the needed MJ moments as well as other diagnostics such as transport coefficients. Measured in the lab frame for an arbitrary distribution, the flux four-vector and stress-energy tensor components are written, respectively, as
$F^{\mu}=\int f(\mathbf{u})u^{\mu}\frac{d^{3}u}{\gamma}$ (5)
$T^{\mu\nu}=m_{0}\int f(\mathbf{u})u^{\mu}u^{\nu}\frac{d^{3}u}{\gamma}.$ (6)
Where the four vector $u^{\mu}$ is defined as $u^{\mu}\equiv(\gamma,u^{i})$. For a comprehensive source on relativistic thermodynamics, see Vereshchagin & Aksenov (2017).
To compute the boost between the lab and fluid-stationary frames we require the velocity $\mathbf{v}_{b}$. The velocity can be computed from the lab frame flux four-vector. Suggestively relabeling some quantities of the flux vector makes this more transparent. First, the component $F^{\mu=0}$ is the density in the lab frame, $N=F^{0}$. Second, for clarity, the fluxes are relabeled $Nv_{b}^{i}=F^{i}$ for $i=\{1,2,3\}$. It is clear then that the velocity is given by $\mathbf{v}_{b}=F^{i}/F^{0}$.
The velocity, $\mathbf{v}_{b}$, gives us all the information we need to transform the lab frame quantities to the fluid-stationary frame. Computing $\Gamma=1/\sqrt{1-v_{b}^{2}}$, where $v_{b}^{2}=\mathbf{v}_{b}\cdot\mathbf{v}_{b}$, the components of the Lorentz transform in Cartesian coordinates is given by
@@
We will call this specific Lorentz transform from the lab to the fluid-stationary frame $\Lambda$. The fluid-stationary frame flux and stress-energy tensor are computed by transforming the lab respective quantities,
$F^{\prime\alpha}=\Lambda_{\phantom{\mu}\mu}^{\alpha}F^{\mu}$ (8)
$T^{\prime\alpha\beta}=\Lambda_{\phantom{\mu}\mu}^{\alpha}\Lambda_{\phantom{\nu}\nu}^{\beta}T^{\mu\nu}.$ (9)
The transformed four-vector for the flux, $F^{\prime\alpha}=(n,0,0,0)$, directly provides the fluid-stationary frame density $n$. Equipped with $T^{\prime\alpha\beta}$, $F^{\alpha}$, and $F^{\prime\alpha}$, we can immediately compute the fluid-stationary frame density, velocity, and temperature,
$n=F^{\prime 0}=N/\Gamma$ (10)
$v_{b}^{i}=F^{i}/F^{0}$ (11)
$P=nT=\frac{1}{3}(T^{\prime 11}+T^{\prime 22}+T^{\prime 33}).$ (12)
Since we only require the diagonal spatial components, we can simplify the expression for the fluid-stationary frame pressure moment calculation to $\mu=\nu=1,2,3$, which yields,
$P=nT=\frac{m_{0}}{3}\int f^{\prime}(\mathbf{u}^{\prime})u^{\prime 2}\frac{d^{3}u^{\prime}}{\gamma^{\prime}}.$ (13)
We can replace the fluid-stationary quantity, $u^{\prime 2}$, in the integrand with lab frame quantities, $u^{\prime 2}=\gamma^{\prime 2}-1=\Gamma^{2}(\gamma-\mathbf{v}_{b}\cdot\mathbf{u})^{2}-1$ from equation (3). In addition, the Lorentz transformation between the primed and unprimed volumes are
@@
The last line, equation (16), is a statement that the phase-space volume element is invariant. These results, pointed out by Zenitani (2015), are arrived at by considering proper or canonical intervals of time, length, and momentum. For example, in the case of time, $d\tau$ = $dt/\gamma$ = $dt^{\prime}/\gamma^{\prime}$, where $\tau$ is the proper/canonical time interval. Therefore, transforming the volume element with equation (14), and using $f(\mathbf{u})=f^{\prime}(\mathbf{u^{\prime}})$, we have an expression for the fluid-stationary pressure computed with the lab frame quantities
$P=nT=\frac{m_{0}}{3}\int f(\mathbf{u})(\Gamma^{2}(\gamma-\mathbf{v}_{b}\cdot\mathbf{u})^{2}-1)\frac{d^{3}u}{\gamma}.$ (17)
Dividing out the fluid-stationary frame number density, $n$, leaves us with the temperature, $T$, of the distribution in the fluid-stationary frame. For general d-dimensional momentum space, the $1/3$ coefficient appearing in equations (12), (13) and (17) becomes $1/d$. In these cases, the integrals and summation are only taken over d-dimensional momentum space.
In summary, the necessary equations to calculate the moments begin with equations (5) and (11) to isolate $\mathbf{v}_{b}$. Then use $\mathbf{v}_{b}$ to compute $\Gamma=1/\sqrt{1-v_{b}^{2}}$ and then to compute $n$ via equation (10). Finally, $T$ is computed using equation (17). This provides all the moments of an arbitrary distribution needed to construct a Maxwell-Jüttner distribution with equivalent moments.
@@ 3.1 Division and Multiplication in DG, Computing $\Gamma$
DG schemes have two equivalent representations of the local solution. One is a modal representation, in which functions are expanded in terms of a finite basis set within each cell, $\phi_{i}$. The second is a nodal representation, in which functions are expanded in terms of a series of points within a cell and multiplied by the associated interpolating polynomials. This choice has tradeoffs, but in this section and the next we choose to consistently use a polynomial order $p$ modal representation for the DG scheme. This is because in previous studies we have found that we can utilize modal representations and corresponding weak operations to minimize aliasing errors, which can greatly impact robustness in kinetic calculations by exciting spurious oscillations. See Juno et al. (2018), Hakim & Juno (2020) and Hakim et al. (2020). We note that it may be favorable to use nodal representations and tolerate the aliasing errors for these computations because, even if they are more inaccurate, the procedure may be more robust to realizability issues.
Choosing the polynomial order of the scheme involves a tradeoff. The benefits of higher polynomial orders are increased accuracy, reduced number of required cells, as well as lower numerical diffusion. However, higher polynomial order expansions increase the computational cost and required memory due to the larger number of degrees of freedom in the basis expansion. In higher dimensional simulations, where the phase space can be 4 to 6 dimensional, the number of basis functions grows rapidly. For instance, if a problem has $d$ dimensions and $p+1$ basis functions per dimension, then the total number of basis functions given by the tensor product is $(p+1)^{d}$. In the simulations presented in a later section using the Gkeyll code, we have found from empirical testing that $p=2$ is the optimal polynomial order for solving the kinetic equation.
The linear algebra calculations, which arise from weak equality of DG-represented values, can be solved exactly. Take, for example, the calculation of isolating $\mathbf{v}_{b}$ from $\langle N\mathbf{v}_{b}\rangle$. Here, the braces, $\langle$$\rangle$, denotes a combined quantity resulting from the moment computation. The expansion of the moments $N,\langle N\mathbf{v}_{b}\rangle$ in terms of an orthonormal set of basis functions in the configuration space, $\phi_{i}(\mathbf{x})$, where $\mathbf{x}$ is defined for an interval within a cell $x_{i}\in[-1,1]$ that has an associated volume, $I$. In each cell, DG quantities are represented as the sum of coefficients times the modal basis
@@
The basis functions $\phi_{i}(\mathbf{x})$ lie in a finite-dimensional function space of order $p$. One such example is the Serendipity basis, presented by Arnold & Awanou (2011), which generalizes the Legendre basis to multiple dimensions while removing interior degrees of freedom to reduce the total number of functions required to represent the solution within each cell. While we use the Serendipity basis to represent the solutions in the simulations presented in later sections, the algorithm does not depend on the choice of basis. The only requirement is that an inner product structure exists to isolate coefficients, for example, in computing the coefficients of $\mathbf{v}_{b}$ from $\langle N\mathbf{v}_{b}\rangle$ and $N$.
With expressions (18), we can rewrite $N\mathbf{v}_{b}=\langle N\mathbf{v}_{b}\rangle$ in terms of their expansions. Where the left-hand side is two separate quantities $N$ times $\mathbf{v}_{b}$ and the right-hand side is a single quantity $\langle N\mathbf{v}_{b}\rangle$. Integrating both sides of this equation gives a relation for the coefficients
$\sum_{i,j}\left(\int_{I}\phi_{i}\phi_{j}\phi_{k}d\mathbf{x}\right)c_{N,i}\mathbf{c}_{v_{b},j}=\sum_{i}\left(\int_{I}\phi_{i}\phi_{k}d\mathbf{x}\right)\mathbf{c}_{Nv_{b},i}.$ (19)
Evaluating the integrals in parenthesis, the right hand side simplifies to the Kronecker delta, $\delta_{ik}$, for an orthonormal basis. The left hand side integral is a bit more involved, but can be calculated analytically before the simulation begins. The weak equality, equation (19) then can be written as
@@
where $M_{ijk}=\int_{I}\phi_{i}\phi_{j}\phi_{k}d\mathbf{x}$ is a 3D array contracted over the $i^{th}$ index with $c_{N,i}$. We can relabel the result as a 2D matrix: $A_{jk}=\sum_{i}M_{ijk}c_{N,i}$. Now this is a clear linear algebra problem. We then find the inversion of $A_{jk}$ to isolate the coefficients $\mathbf{c}_{v_{b},j}$ which solves the linear system and isolates $\mathbf{v}_{b}$
@@
Hypothetically, if we desired to calculate $\mathbf{c}_{Nv_{b},k}$ given the other two coefficients, $c_{N,i}$, $\mathbf{c}_{v_{b},j}$ we could have simply stopped at equation (20).
@@
$c_{\Gamma,i}=\int_{I}\frac{1}{\sqrt{1-\sum_{m}\left(\mathbf{c}_{v_{b},m}\phi_{m}(x)\right)^{2}}}\phi_{i}(x)d\mathbf{x}\approx\sum_{k}\frac{w_{k}\phi_{i}(x_{k})}{\sqrt{1-\sum_{m}\left(\mathbf{c}_{v_{b},m}\phi_{m}(x_{k})\right)^{2}}}.$ (22)
@@
The approximations in computing the coefficients $c_{\Gamma,i}$ arise from projecting the nonlinear result back onto the finite polynomial basis, even when the result contains higher order terms. Operations, such as multiplication, square roots, and exponents of quantities represented on the grid will likewise result in polynomial orders greater than the basis. Projecting the result of these operations back onto the basis gives us a weakly equal projection, where the result and its projection are not pointwise equal, but have matching basis coefficients.